OpenStudy (anonymous):

Find the slope of the line passing through (4, 9) and (-2, 9). a. 0 b. undefined

OpenStudy (anonymous):

a

OpenStudy (anonymous):

a

OpenStudy (mertsj):

slope is rise/run So subtract the y values and subtract the x values. Put the y difference over the x difference and simplify. That is your slope. 9-9=0 4-(-2)=6 0/6=0